# Multivariate Diffusion Transformer with Decoupled Attention for

### v1

@@ Appendix A Supplementary Material
Comparisons TediGAN (xia2021tedigan) GAN 0.6707 0.7257 2.367 71.55 22.44 5.50 -0.2374
UaC (nair2023unite) Diffusion 0.3647 0.6983 3.636 74.33 22.38 4.66 -1.2595
PixelFace+ (du2023pixelface+) GAN 0.3214 0.6987 3.304 91.20 23.91 4.65 -1.2848
CoDiffusion (huang2023collaborative) Diffusion 0.3825 0.7556 2.221 53.44 21.83 5.85 1.4839
ControlNeXt (peng2024controlnext) Diffusion 0.5483 0.7576 1.138 74.14 27.05 5.12 -0.5563
MM2Latent (meng2025mm2latent) GAN 0.7552 0.7179 1.390 67.43 24.05 5.21 -0.0827
@@ A.1 Additional Quantitative Experiments
Table 7 presents the quantitative evaluation results of our proposed MDiTFace method for multimodal facial synthesis, with experiments carried out on the MM-FairFace dataset. Built on the FairFace dataset (karkkainen2021fairface) (released by MIT in 2019 to reduce race-, gender-, and age-related biases in face recognition), MM-FairFace inherits race, gender, and age annotations for each facial image. We constructed MM-FairFace by using a pre-trained facial parser (zheng2022general) to generate semantic masks for each image, following a methodology similar to that for the MM-FFHQ dataset.
@@ A.2 Qualitative Comparisons with SOTA Methods
Figures 7, 8, and 9 display visual comparisons of typical facial images synthesized by our proposed MDiTFace method and competing state-of-the-art approaches on the MM-CelebA (lee2020maskgan), MM-FFHQ (zhou2021generative), and MM-FairFace (karkkainen2021fairface) datasets, respectively. Visually, our method yields facial images with superior naturalness and realism, accompanied by a marked reduction in artifacts. Moreover, compared to other methods, MDiTFace shows distinct advantages in semantic mask alignment (reflecting spatial layout) and fine-grained attribute matching (e.g., hair color, accessories, facial makeup).
Notably, on the MM-FairFace dataset (karkkainen2021fairface), many competing methods exhibit facial synthesis collapse due to their failure to interpret racial annotations, resulting in poor generalization. In contrast, MDiTFace shows superior robustness by rationally deducing skin tone from textual racial cues, highlighting its strong generalization in textual semantic comprehension.
@@ A.4 Diversity in Facial Generation
Clip2Latent (pinkneyclip2latent) 0.7861 1.410 27.56 1.0129
GCDP (park2023learning) 0.6118 1.456 25.94 0.4563
E3-FaceNet (zhang2024fast) 0.7319 2.749 27.94 0.8150
@@
INADE (tan2021diverse) 0.6156 1.871 93.60 2.57
E2Style (wei2022e2style) 0.8007 1.129 90.68 2.36
SemFlow+ (wang2024semflow) 0.7098 1.767 94.35 2.30

### v2

@@ Appendix A Supplementary Material
Comparisons TediGAN (2021) GAN 0.6707 0.7257 2.367 71.55 22.44 5.50 -0.2374
UaC (2023) Diffusion 0.3647 0.6983 3.636 74.33 22.38 4.66 -1.2595
PixelFace+ (2023) GAN 0.3214 0.6987 3.304 91.20 23.91 4.65 -1.2848
CoDiffusion (2023) Diffusion 0.3825 0.7556 2.221 53.44 21.83 5.85 1.4839
ControlNeXt (2024) Diffusion 0.5483 0.7576 1.138 74.14 27.05 5.12 -0.5563
MM2Latent (2025) GAN 0.7552 0.7179 1.390 67.43 24.05 5.21 -0.0827
@@ A.1 Additional Quantitative Experiments
Table 7 presents the quantitative evaluation results of our proposed MDiTFace method for multimodal facial synthesis, with experiments carried out on the MM-FairFace dataset. Built on the FairFace dataset (Karkkainen and Joo 2021) (released by MIT in 2019 to reduce race-, gender-, and age-related biases in face recognition), MM-FairFace inherits race, gender, and age annotations for each facial image. We constructed MM-FairFace by using a pre-trained facial parser (Zheng et al. 2022) to generate semantic masks for each image, following a methodology similar to that for the MM-FFHQ dataset.
@@ A.2 Qualitative Comparisons with SOTA Methods
Figures 7, 8, and 9 display visual comparisons of typical facial images synthesized by our proposed MDiTFace method and competing state-of-the-art approaches on the MM-CelebA (Lee et al. 2020), MM-FFHQ (Zhou 2021), and MM-FairFace (Karkkainen and Joo 2021) datasets, respectively. Visually, our method yields facial images with superior naturalness and realism, accompanied by a marked reduction in artifacts. Moreover, compared to other methods, MDiTFace shows distinct advantages in semantic mask alignment (reflecting spatial layout) and fine-grained attribute matching (e.g., hair color, accessories, facial makeup).
Notably, on the MM-FairFace dataset (Karkkainen and Joo 2021), many competing methods exhibit facial synthesis collapse due to their failure to interpret racial annotations, resulting in poor generalization. In contrast, MDiTFace shows superior robustness by rationally deducing skin tone from textual racial cues, highlighting its strong generalization in textual semantic comprehension.
@@ A.4 Diversity in Facial Generation
Clip2Latent (2022) 0.7861 1.410 27.56 1.0129
GCDP (2023) 0.6118 1.456 25.94 0.4563
E3-FaceNet (2024) 0.7319 2.749 27.94 0.8150
@@
INADE (2021) 0.6156 1.871 93.60 2.57
E2Style (2022) 0.8007 1.129 90.68 2.36
SemFlow+ (2024) 0.7098 1.767 94.35 2.30